Atomic weight of Ne  is 20.02 . Ne is mixture of Ne20  and Ne22 relative , abundance of heavier isotope is ;
 
(1)  90       (2) 20
(3)  40        (4) 10

Grade:12th pass

1 Answers

Mohamed shakeel
5 years ago
there is some discrepancy in the question. If the atomic weight of neon given in the question is 20.2 then your answer is option 4]10.
please check it because
 20x+22[100-x]/100  =20.2
20x+2200-22x =2020
180-2x =0
2x=180
x =90 %
if x= 90%
then abundance of heaver isotope is 100-x =100-90= 10%
 
if 20.02 then
20x+22[100-x]/100 =20.02
20x+2200-22x =2002
198-2x =0
then on solving for x
x= 99%
then abundance of heavier isotope is 100-99=1%
please check your question if the weight given is 20.2 or 20.02
 
please approve my answer